### v2

# Learning to Balance: Diverse Normalization for Cloth-Changing Person Re-Identification

Abstract
Cloth-Changing Person Re-Identification (CC-ReID) involves recognizing individuals in images regardless of clothing status. In this paper, we empirically and experimentally demonstrate that completely eliminating or fully retaining clothing features is detrimental to the task. Existing work, either relying on clothing labels, silhouettes, or other auxiliary data, fundamentally aim to balance the learning of clothing and identity features. However, we practically find that achieving this balance is challenging and nuanced. In this study, we introduce a novel module called Diverse Norm, which expands personal features into orthogonal spaces and employs channel attention to separate clothing and identity features. A sample re-weighting optimization strategy is also introduced to guarantee the opposite optimization direction. Diverse Norm presents a simple yet effective approach that does not require additional data. Furthermore, Diverse Norm can be seamlessly integrated ResNet50 and significantly outperforms the state-of-the-art methods.
Introduction
Person Re-identification (ReID), which aims to match a target person’s image across different camera views, is highlighted as a crucial task in the field of intelligent surveillance systems (Shi, Liu, and Liu 2020) and multi-object tracking (Ke et al. 2019). In the early works of ReID, researchers normally take the assumption that people don’t change their clothes in a short period of time. Therefore, methods developed during this stage mostly leverage the clothing information to identify people (Gu et al. 2020; Zhou et al. 2019a; Zheng et al. 2015; Gu et al. 2019; Hou et al. 2021). Generally, these methods can perform well in short-term datasets but will suffer from significant performance degradations when testing on a long-term one where people change their clothes frequently. To overcome such limitations, researchers has gradually turn their attention to consider the change of clothes in model training and testing.
Figure 1: A trade-off exists between maintaining clothing consistency and clothing changes features in CC-ReID. Traditionally, the model must acquire two distinct sets of features: one focused on clothing features (such as garments and trousers) and another on clothing-irrelevant attributes (like face and body shape) to effectively handle gallery with same person involving cloth consistency and cloth changing, respectively. Finally, either clothes or ID-invariant features will be used to match galleries images of the same person with different clothing status.
With years of developments, the mainstream solutions of cloth-changing person re-identification (CC-ReID) are to either incorporate more ID features from other modalities (e.g., contour sketch (Yang, Wu, and Zheng 2019), body shape (Qian et al. 2020a), hairstyle (Wan et al. 2020a; Yu et al. 2020) and 3D shape (Chen et al. 2021a)) and encourage the model to learn from them, or use hand-crafted clothes labels (Gu et al. 2022; Cui et al. 2023; Yang et al. 2023) to force the model pay less attention to the clothes appearance in CC-ReID task. However, as illustrated in Fig. 1, regardless of which path is taken, we argue that all these works are actually trying to find a best trade-off point for encoding ID features and clothing features into the high-level representation of the model. To further validate this conjecture, we conducted grid searching experiments on CAL (Gu et al. 2022) and illustrated the result in Fig. 2. Specifically, we adjusted the intensity of the adversarial loss, denoted as $\epsilon$. A larger $\epsilon$ indicates the removal of more clothing features. We found that completely removing clothing features is detrimental; it can also negatively impact scenarios where clothing remains unchanged. Conversely, retaining too many clothing features is detrimental to scenarios involving clothing changes. The reason for this phenomenon is quite intuitive. As illustrated in Fig. 1, both the query and gallery datasets in CC-ReID contain scenarios of both clothes-changing and clothes-consistency person. Therefore, the optimal model represents the best equilibrium point between these two scenarios.
Figure 2: The relation between accuracy and strength of removing clothes features in CAL (Gu et al. 2022). We found that completely removing clothing features is not beneficial; instead, it can simultaneously disrupt scenarios where clothing is not changed. However, if too many clothes features are retained, it will not be good for the clothes changing scenes.
When viewing the CC-ReID problem from the “trade-off” perspective, we find that the previous efforts of using multi-modality or clothes labels are all for separating the learning of clothing and ID features so as to balance their proportion. However, we notice that it is unnecessary with so much effort because the learnings of clothing and ID features are naturally separated to each other (i.e., a feature cannot simultaneously belongs to both clothing-relevant and ID-invariant feature groups). In Fig. 2, we further notice that “trade-off” searching in existing works is actually an ill-posed optimization problem with two optimization directions fighting with each other in the same model. Therefore, we can actually separating such learning using only RGB modality and free from acquiring clothes labels with only a simple whitening operation. Inspired by the ideas from ensemble learning, this paper explicitly constructs two separate branches with each one focusing on one optimization target, and the diversity between them is ensured by a whitening operation (Huang et al. 2018b) and channel attention (Hu, Shen, and Sun 2018). Moreover, sampling re-weighting strategy (Kim et al. 2022) is also used to serve as a regularization to assign opposite sample weights to let different branches focus on specific input samples. For instance, the back-view sample is heavily weighted towards the clothing branch, while the face-forward sample is prioritized for the ID branch.
Contributions: 1) For the first time, we points out that a potential bottleneck in existing work may arise from ill-posed optimization issue and a tendency to become trapped in local minima. 2) To address this issue, we present Diverse Norm for CC-ReID, which is designed to create orthogonal features from the same person embeddings, allowing for meeting the requirement of short-term and long-term ReID. 3) Our provide compelling evidence that Diverse Norm surpasses other state-of-the-art methods, including both multi-modality-based and clothing labels-based approaches.
Figure 3: Architecture Overview of Our Method. In Fig. (a), we first apply whitening to the features extracted by the backbone network and then utilize channel attention to separate clothing and identity features. To effectively distinguish between clothes and identity features, we constructed two classifiers and employed sample reweighting to achieve the concept selection. Specifically, as shown in Fig. (b), when an individual is in a dimly lit area wearing dark clothing, making recognition based solely on clothing difficult, the identity branch increases its weight. Conversely, as depicted in Fig. (c), when a person is moving quickly, causing facial motion blur, but wearing distinct plaid clothing, the apparel branch increases its weight.
Related Work
Long-Term Person Re-Identification. The primary focus in clothes-changing re-identification (re-id) research is to extract features from images that are not influenced by clothing. Within this field, two main approaches have emerged: 1) A branch of research attempts to use disentangled representation learning (Zhang et al. 2019; Zheng et al. 2019; Chan et al. 2023; Gu et al. 2022) to separate appearance from structural information in RGB images. This approach treats structural information as features that are independent of clothing. 2) Another branch of research aim to combine multi-modality data, such as, skeletons (Feng, Li, and Luo 2016; Ariyanto and Nixon 2012), silhouettes (Chao et al. 2019; Han and Bhanu 2006), 3D information (Liu et al. 2023; Chen et al. 2021b) to exert the structural information. However, we found that achieving this balance is fragile. Therefore, this paper attempts to optimize the CC-ReID problem using a decoupled approach.
Whitening and Orthogonality. In deep learning, orthogonality constraints have been applied to address vanishing or exploding gradients, particularly in Recurrent Neural Networks (RNNs) (Vorontsov et al. 2017; Mhammedi et al. 2017; Wisdom et al. 2016). These constraints have been extended to various neural network types, including non-RNNs (Harandi and Fernando 2016; Huang et al. 2018a; Lezcano-Casado and Martínez-Rubio 2019; Lezama et al. 2018). Some methods use specialized loss functions to enforce orthogonality (Lezama et al. 2018). In contrast, CW optimizes the orthogonal matrix using Cayley-transform-based curvilinear search algorithms (Wen and Yin 2013), with the unique goal of aligning orthogonal matrix columns with specific concepts, distinguishing it from prior methods. Our Diverse Norm module incorporates whitening techniques from IterNorm (Huang et al. 2019a) to separate appearance from structural information in RGB images.
Methodology
Diversity Normalization. As mentioned in Fig. 2, achieving a balance between clothing and identity features is challenging, and even slight missteps can degrade the performance of one aspect. To address this issue, we introduce an expand operation, aiming to learn robust representations of both clothing and identity features during training. Specifically, we introduce whitening (Huang et al. 2018b) to disentangle the conception (Chen, Bei, and Rudin 2020), combined with channel attention (CA) (Hu, Shen, and Sun 2018) for clothes and identity features selection. Formally, whitening is defined as:
Definition 1
Whitening is designed to process latent features in a latent space through decorrelation, standardization (whitening), and orthogonal transformation. The whitening transformation can be represented as:
$\psi(\mathbf{Z})=\mathbf{W}(\mathbf{Z}-\mathbf{\mu}),$ (1)
where $\mathbf{\mu}$ is the sample mean and $\mathbf{W}$ is the whitening matrix satisfying $\mathbf{W}^{\top}\mathbf{W}=\mathbf{\Sigma}^{-1}$, with $\mathbf{\Sigma}$ is the covariance matrix.
Given the human features $h$ generated by the backbone network, we then integrate CA into our architecture to separate clothing and identity features. Specifically, the channel weights in the CA block are calculated as:
$\omega=\sigma\left(Wg(\psi(h))\right)\in(0,1),$
where $g$ is the channel-wise Global Average Pooling (GAP), $W$ is fully-connected layer, and $\sigma$ is the Sigmoid function. As clothing and identity features do not spatially overlap, we use $\omega$ and $1-\omega$ to decompose $\hat{h}$ into $h_{ID}$ and $h_{C}$:
$h_{id}=\psi(h)\cdot\omega,\quad h_{c}=\psi(h)\cdot(1-\omega)$
where $h_{id}$ and $h_{c}$ represent identity features and clothing features, respectively.
Sample Re-weighting for Conception Separation. Obviously, the whitening and CA only guarantee the orthogonality of feature pairs and feature selection, which necessitates further strategies to ensure the separation between clothing-relevant and clothing-irrelevant features. As shown in Fig. 3, we observe that among training samples, some are identifiable by clothing, while others are recognizable by identity features. Based on this observation, we employ a re-weighting strategy (Nam et al. 2020; Creager, Jacobsen, and Zemel 2021) to amplify this distinction, encouraging classifiers to be biased towards certain types of samples. Formally, given the classifier output $\tilde{y}(h_{id})$ and $\tilde{y}(h_{c})$, the relative difficulty score $\mathcal{W}(h_{id})$ and $\mathcal{W}(h_{c})$ are defined as follows:
$\mathcal{W}(h_{id})\equiv 1,\ \mathcal{W}(h_{c})=\frac{2\mathcal{L}_{ID}\left(\tilde{y}(h_{c}),y\right)}{\mathcal{L}_{ID}\left(\tilde{y}(h_{id}),y\right)+\mathcal{L}_{ID}\left(\tilde{y}(h_{c}),y\right)},$
where $\mathcal{W}(h_{id})$ serves as an anchor (Creager, Jacobsen, and Zemel 2021), and always equals 1, $\mathcal{W}(h_{c})$ is the ratio of loss between $h_{id}$ and $h_{c}$, and $\mathcal{L}_{ID}$ indicates cross entropy loss for classification. We draw upon (Creager, Jacobsen, and Zemel 2021) to define the anchor, which represents a biased branch. In 2, the baseline is assumed to be biased towards the branch that predominantly captures clothing features. This assumption is reasonable, considering that clothing can account for up to 50% of the visual content in an image. Hence, the relative difficulty score $\mathcal{W}(h_{c})$ is employed to guide the identity classifier during its learning process.
Similarity between Query and Gallery Data. Let $q_{id}$, $q_{c}$, $g_{id}$, and $g_{c}$ represent the identity and clothing features of the query and gallery images, respectively. We employ cosine distance to measure the similarity between the feature of the query and gallery images and then aggregate these similarities. Our experiments indicate that this approach significantly enhances model performance. We also compared the effects of combining $q_{id}+q_{c}$ and $g_{id}+g_{c}$ in Fig. 6. We found that the primary reason for the traditional baseline’s failure is the overshadowing of ID features, which account for a much smaller proportion compared to clothing features.
Table 1: Comparison with state-of-the-art methods on LTCC, and PRCC. The results indicated by an underline were copied from the original papers. A backslash indicates that a result is not available. The Extra data indicates, such as, contour sketches, silhouettes, human poses, and 3D shape information.
Methods Cloth labels Extra data LTCC PRCC
General CC Same Clothes CC
mAP Rank1 Rank5 mAP Rank1 Rank5 mAP Rank1 Rank5 mAP Rank1 Rank5
[…]
3APF $\checkmark$ 21.37 54.41 70.23 9.72 22.43 38.76 95.02 98.04 99.69 41.63 43.95 54.58
[…]
CASE $\checkmark$ 21.10 52.70 70.59 9.86 22.96 38.52 96.18 98.89 99.41 36.68 42.90 56.34
FD-GAN $\checkmark$ 36.89 73.43 80.12 15.36 32.91 46.68 99.69 99.97 100.00 58.57 58.34 63.82
[…]
ReIDCaps 15.81 43.03 58.61 6.20 12.82 23.71 96.62 99.54 99.61 44.81 48.00 54.61
BoT 27.80 64.33 76.91 10.70 28.12 42.12 98.41 99.30 99.82 45.31 44.92 52.51
PCB 30.39 64.50 76.27 11.24 21.27 38.78 97.01 99.02 99.90 37.56 36.49 43.75
Part-aligned 21.33 53.18 69.61 9.58 21.41 38.50 95.63 98.70 99.82 39.33 42.24 51.51
MGN 33.85 66.32 74.95 13.69 27.30 41.34 98.77 99.92 99.97 53.12 55.21 62.09
SCPNet 18.02 48.07 67.34 7.38 15.82 31.63 93.09 97.68 99.10 23.48 22.92 35.39
ISGAN 33.76 68.56 77.69 $\underline{13.03}$ 29.08 42.86 99.65 99.90 100.00 55.97 55.38 61.84
DGNet 32.18 64.71 78.30 10.80 21.70 34.69 97.42 99.15 99.85 48.17 49.08 63.53
ResNet50 29.41 65.57 73.80 11.05 28.16 42.13 98.44 97.90 99.81 43.36 45.67 52.32
+Diverse Norm 46.48 78.92 82.45 31.88 63.30 71.97 99.89 99.94 100.00 58.87 59.74 64.42
TransReID 33.08 63.74 74.00 15.17 26.88 43.19 99.89 99.99 100.00 43.08 48.96 51.22
+Diverse Norm 47.08 77.75 83.99 33.38 60.70 72.33 99.89 99.91 100.00 59.28 61.35 67.98
Table 2: Comparison with state-of-the-arts on VC-Clothes
Method General SC CC
(all cams) (cam2&cam3) (cam3&cam4)
top-1 mAP top-1 mAP top-1 mAP
MDLA 88.9 76.8 94.3 93.9 59.2 60.8
PCB 87.7 74.6 94.7 94.3 62.0 62.2
Part-aligned 90.5 79.7 93.9 93.4 69.4 67.3
FSAM - - 94.7 94.8 78.6 78.9
3DSL - - - - 79.9 81.2
CAL 92.9 87.2 95.1 95.3 81.4 81.7
ResNet50 88.3 79.2 94.1 94.3 67.3 67.9
+Diverse Norm 93.1 88.2 96.6 95.8 85.9 83.2
Table 3: Comparison on LaST and DeepChange.
Method LaST DeepChange
top-1 mAP top-1 mAP
OSNet 63.8 20.9 39.7 10.3
ReIDCaps - - 39.5 11.3
BoT 68.3 25.3 47.5 13.0
mAPLoss 69.9 27.6 - -
CAL 73.7 28.8 54.0 19.0
ResNet50 69.4 25.6 50.6 15.9
+Diverse Norm 74.4 29.7 56.9 19.4
Experiment
In this paper, our experiment firstly involves two key CC-ReID datasets, PRCC (Yang, Wu, and Zheng 2021) and LTCC (Qian et al. 2020b), which are essential for the current state of the arts’s comparison. We here aim to prove that the the clothes labels for the CC-ReID is not necessary if we explicitly disentangled the clothes and identical features. PRCC dataset has 22,889 training and 10,800 test images, with test samples from A forming the gallery set, and B and C forming the query set for different matching scenarios (Yang, Wu, and Zheng 2021). LTCC, similar to PRCC, comprises 9,576 training images and 7,050 test images. The dataset is divided for same-clothes and cross-clothes matching, with 493 randomly selected query images from various cameras and outfits (Qian et al. 2020b). In LTCC, clothing information is readily available, while in PRCC, the clothing labels can be inferred conveniently from the camera IDs, which are used in 3DSL (Chen et al. 2021b), CASE (Li, Weng, and Kitani 2021), FD-GAN (Chan et al. 2023), CAL (Gu et al. 2022), CVSL (Nguyen et al. 2024).
During the evaluation phase, ResNet50 with Diverse Norm calculates the average similarity of branches $h_{id}$ and $h_{c}$ between each query image and gallery image for each person image in the query set, respectively. The performance of the ReID models is assessed using mean Average Precision (mAP) and Cumulative Matching Characteristics (CMC) at Rank1 and Rank5 matching accuracy. In the case of PRCC, CMC results are provided for our model under both single-shot and multi-shot settings.
Figure 4: Comparing the effects of training ResNet50 with and without Diverse Norm on the LTCC dataset.
Implementation Details. We use ResNet50 (He et al. 2016) as the backbone architecture, omitting the final downsampling step for finer granularity. For LTCC and PRCC, following prior work (Gu et al. 2022), we apply global average pooling and global max pooling to the output feature map, concatenate the pooled features, and use task-specific BatchNorm (Ioffe and Szegedy 2015) for normalization. Following (Qian et al. 2020b), input images are resized to $384\times 192$. Data augmentation includes random horizontal flipping, cropping, and erasing (Zhong et al. 2020). The batch size is 64, with 8 individuals per batch, each represented by 8 images. We train the model with the Adam optimizer (Kingma and Ba 2015) for 60 epochs, using a channel attention module (Hu, Shen, and Sun 2018) as a projection layer before Diverse Norm. The initial learning rate is $3.5e^{-4}$, reduced by a factor of 10 every 20 epochs.
Comparison with State-of-the-art Method
We compared our proposed Diverse Norm with several state-of-the-art ReID models, both for standard ReID and those designed for clothes-changing scenarios. Standard ReID models include BoT (Luo et al. 2019a), PCB (Sun et al. 2018), MGN (Wang et al. 2018), SCPNet (Fan et al. 2019), ISGAN (Eom and Ham 2019), AIM (Yang et al. 2023), CAL (Gu et al. 2022), and OSNet (Zhou et al. 2019a), while clothes-changing ReID models include LTCC (Qian et al. 2020a), CASE (Li, Weng, and Kitani 2021), FSAM (Hong et al. 2021), CCFA (Han et al. 2023) and 3DSL (Chen et al. 2021b). All models are implemented by Chan et al.(Chan et al. 2023) using available codes, except for those without open-source code (such as (Li, Weng, and Kitani 2021; Qian et al. 2020a; Wan et al. 2020a; Yang, Wu, and Zheng 2019)), which were implemented based on paper descriptions.
Figure 5: The connection between model performance and the number of clothes on LTCC.
Evaluation on LTCC and PRCC. The results on clothes-changing ReID datasets, PRCC and LTCC, are presented in Table 1. It’s evident that all models experience a significant drop in accuracy when individuals change their clothes. In scenarios without clothing changes on PRCC, all models perform admirably, with most achieving over 96% in Rank 1. Half of them even exceed 99% in Rank 1, except ASE-SPT, which relies heavily on sketch images and achieves 72.09%. Surprisingly, models designed for clothing changes do not outperform traditional models when clothing changes are absent. These specialized ReID models do not demonstrate a clear advantage over traditional models in the clothing-changing scenario in both datasets. Some traditional models even outperform clothing-changing models. For example, BoT achieves 44.92% in Rank 1 on PRCC with clothing changes and 64.33% and 28.12% in the two settings on LTCC, outperforming most clothing-changing models. MGN and ISGAN even achieve over 50% in Rank 1 on PRCC with clothing changes.
In Table 1, several clothing-changing models, like CASE, 3APF, ReIDCaps, CAL, AIM, and AFD-Net, perform above average but fall short of achieving high results compared to Diverse Norm, especially in LTCC dataset. CASE encourages its identity encoder to overlook color information, potentially missing out on important clothing details other than color. Similarly, ReIDCaps learns clothes information across different dimensions of the output vector, making it challenging to ensure proper disentanglement of clothing features. AFD-Net also extracts and disentangles identity and clohtes features, where the latter is assumed to represent clothing features. Nevertheless, there is no assurance that identity details are successfully acquired and separated in the learning process. In contrast, the Diverse Norm method explicitly orthogonalize identity and clothing features, resulting in a more favorable optimization landscape. Surprisingly, by simply integrating this module into the backbone, we achieve state-of-the-art performance in both scenarios involving consistent clothing and clothing changes.
Figure 6: Comparing different query strategies.
Evaluation on VC-Clothes. To further demonstrate the superior of Diverse Norm, the VC-Clothes dataset (Wan et al. 2020b) is also considered, which is a synthetic virtual dataset generated using GTA5. It comprises a total of 19,060 images depicting 512 different identities captured from 4 distinct camera scenes. Each identity is seen wearing 1 to 3 different outfits, and it’s worth noting that all images of the same identity captured by cameras 2 and 3 feature them wearing identical clothing. Consequently, many recent studies (Wan et al. 2020b; Hong et al. 2021) evaluate performance on a subset of this dataset obtained from cameras 2 and 3, referred to as the same-clothes (SC) setting. Additionally, results are reported for a subset from cameras 3 and 4, referred to as the clothes-changing (CC) setting. In this work, we adhere to these established evaluation settings for the sake of fair comparison. To assess the effectiveness, we benchmark it against two single-modality-based re-id methods MDLA (Qian et al. 2017) and PCB (Sun et al. 2018), one clothes label based method CAL (Gu et al. 2022) as well as three multi-modality-based re-id methods Part-aligned (Suh et al. 2018), FSAM (Hong et al. 2021), and 3DSL (Chen et al. 2021b) on the VC-Clothes dataset. The results are summarized in Table 3. Notably, with two expanding branches, Diverse Norm consistently outperforms both the baseline and all the aforementioned state-of-the-art methods across all evaluation metrics, including general performance, the same-clothes setting, and the clothes-changing setting.
Evaluation on Large-scale Datasets. LaST (Shu et al. 2021) and DeepChange (Xu and Zhu 2021) are two large-scale long-term person re-id datasets. LaST is a large-scale dataset with over 228,000 carefully labeled pedestrian images collected from movies, which is used to study the scenario where pedestrians have a large activity scope and time span. DeepChange is a dataset of 178,000 bounding boxes from 1,100 person identities, collected over 12 months. To ensure a fair comparison (Gu et al. 2022), we set the batch size to 64 and each batch contains 16 persons and 4 images per person. We report the performance of Diverse Norm and other methods: OSNet (Zhou et al. 2019b), ReIDCaps (Huang et al. 2019b), BoT (Luo et al. 2019b), mAPLoss (Shu et al. 2021), and CAL (Gu et al. 2022), in a general setting. On DeepChange, we allow true matches to come from the same camera but different tracklets as query following (Xu and Zhu 2021). As shown in Table 3, Diverse Norm outperforms the baseline and state-of-the-art methods on both LaST and DeepChange. Notably, Diverse Norm remains effective on DeepChange when comparing with CAL, which uses the collection date as a proxy for clothes labels.
Abalation Study
Examining the Impact of Clothes. We observed an interesting phenomenon in Table 1 where the improvement achieved by our method differs between the LTCC and PRCC datasets. We conjecture this discrepancy is due to the most significant difference between LTCC and PRCC: the number of clothes per person. In the PRCC dataset, each individual has only two outfits, whereas in the LTCC dataset, an individual can have up to 14 different outfits. We believe the number of outfits plays a crucial role in the performance of learning features that are unrelated to clothing, thereby affecting the overall performance of our method. To further evaluate our hypothesis, we conducted an experiment by removing different proportions of outfits in the LTCC dataset. The results are shown in Fig. 5. We can observe that as the number of clothes decreases, the model’s performance experiences a significant decline.
Examining Different Query Strategies. In this ablation study, we evaluated two distinct query strategies: one that computes Cosine similarity after summing the features, and another that calculates similarities separately before summation. The results are presented in Fig. 6. Our findings reveal a substantial improvement when employing the separate calculation method. We hypothesize that this improvement stems from the disproportionate representation of clothing in these images, which previously led to a skewed similarity metric that heavily favored clothing features. Due to the extreme imbalance in the proportion of clothing versus facial features within an image, even though the ResNet50 extracts facial features, they are often overshadowed by the dominant clothing features. By calculating Cosine similarities separately, we mitigate this bias, allowing for a more balanced consideration of both clothing and identity-specific features.
Conclusion
In this paper, we are the first to identify the challenges faced by the CC-ReID baseline, specifically its difficulty in balancing the components of clothing and ID features. We introduced a novel technique called Diverse Norm for CC-ReID, which does not require any additional data, multi-modality inputs, or clothing labels. This method simply facilitates the separation of person features into an orthogonal space, effectively distinguishing between clothing and clothing-irrelevant attributes. Notably, Diverse Norm provides a straightforward and highly effective solution that can be seamlessly integrated into any CC-ReID dataset, surpassing current state-of-the-art methods that rely on additional information. We hope that Diverse Norm will become a commonly used module in future clothes-changing person re-identification methods.